# Functional inequalities on the biased and restricted cube: an induction approach

### v2

@@ 2.2 Induction Part
  • We may assume without loss of generality that $g\geq 0$, since replacing $g$ by its absolute value preserves the right-hand side of (5), while it only decreases the left-hand side:
@@
  • We may assume the support of $g$ is the whole set $A$. Indeed, if we denote $B:={\rm supp}(g)=\{x\in\{0,1\}^{n}:g(x)>0\}$ with $B\subset A$. Then $\mathbb{E}_{p}[g]^{2}=\left(\sum_{x\in B}g(x)\mu_{p}(x)\right)^{2}$. Hence, replacing the support of $g$ with the whole set $A$ only increases the right-hand side of (5).
@@
$\begin{split}&p\mathcal{E}_{p}^{n-1}(g_{1},g_{1})\geq\frac{\mathbb{E}_{p}[g_{1}]^{2}}{a_{1}}\log_{p}a_{1}\\ &p\mathcal{E}_{p}^{n-1}(g_{0},g_{0})\geq\frac{\mathbb{E}_{p}[g_{0}]^{2}}{a_{0}}\log_{p}a_{0}.\end{split}$ (15)
@@
    where the first inequality is due to $a_{j}=\mu_{p}^{n-1}(A_{j})\leq 1$ and the second inequality is due to Jensen’s inequality. Hence replacing $g_{j}$ with their averages on the corresponding subsets can only decrease the left-hand side and does not change the right-hand side in the second inequality of (16).
@@
$\begin{split}&s_{1}:=\mathbb{E}_{p}[g_{1}]=\sum\limits_{x\in\{0,1\}^{n-1}}\mathbbm{1}_{A_{1}}(x)\cdot g_{1}(x)\mu_{p}^{n-1}(x)\\ &\quad\ =\sum\limits_{x\in A_{0}}\mathbbm{1}_{A_{1}}(x)\cdot g_{1}(x)\mu_{p}^{n-1}(x)+\sum\limits_{x\in A_{1}\setminus A_{0}}\mathbbm{1}_{A_{1}}(x)\cdot g_{1}(x)\mu_{p}^{n-1}(x)=\alpha a_{0}+\gamma(a_{1}-a_{0}),\\ &s_{0}:=\mathbb{E}_{p}[g_{0}]=\sum\limits_{x\in\{0,1\}^{n-1}}\mathbbm{1}_{A_{0}}(x)\cdot g_{0}(x)\mu_{p}^{n-1}(x)\\ &\quad\ =\sum\limits_{x\in A_{0}}\mathbbm{1}_{A_{0}}(x)\cdot g_{0}(x)\mu_{p}^{n-1}(x)+\sum\limits_{x\in A_{1}\setminus A_{0}}\mathbbm{1}_{A_{0}}(x)\cdot g_{0}(x)\mu_{p}^{n-1}(x)=\beta a_{0},\\ &\|g_{1}-g_{0}\|_{2,\mu_{p}}^{2}=\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}[(g_{1}(x^{\prime})-g_{0}(x^{\prime}))^{2}]\\ &\quad\ =\sum\limits_{x^{\prime}\in A_{0}}(g_{1}(x^{\prime})-g_{0}(x^{\prime}))^{2}\mu_{p}^{n-1}(x^{\prime})+\sum\limits_{x^{\prime}\in A_{1}\setminus A_{0}}(g_{1}(x^{\prime})-g_{0}(x^{\prime}))^{2}\mu_{p}^{n-1}(x^{\prime})\\ &\quad\ =(\alpha-\beta)^{2}\cdot a_{0}+(a_{1}-a_{0})\cdot\gamma^{2}.\end{split}$
@@
$ps_{1}^{2}f(a_{1})+(1-p)s_{0}^{2}f(a_{0})+pa_{0}(\alpha-\beta)^{2}+p(a_{1}-a_{0})\gamma^{2}\geq(ps_{1}+(1-p)s_{0})^{2}f(pa_{1}+(1-p)a_{0})).$ (18)
@@ 2.3 Two-point Inequality Part
Renaming the variables $x=a_{0}$ and $y=a_{1}$, and recalling the constraints on $a_{0}$ and $a_{1}$, we need to prove (20) for $0<x<y\leq 1$.
$pf(y)+(1-p)f(x)+(1-p)yf(x)f(y)\geq(p(1-p)yf(x)+(1-p)^{2}yf(y)+1)f(py+(1-p)x)$ (21)
@@
$\begin{split}\Delta_{p}(x,y)=\Delta_{p}(ty,y)&=pf(y)+(1-p)f(t\cdot y)-f((p+(1-p)t)\cdot y)\\ &=pf(y)+\frac{1}{t}\cdot(1-p)f(y)+\frac{1}{y}\cdot(1-p)f(t)-\frac{f(y)}{p+(1-p)t}-\frac{f(p+(1-p)t)}{y}\\ &=\frac{p(1-p)(t-1)^{2}}{t(p+(1-p)t)}\cdot f(y)+\frac{1}{y}\cdot\left(\frac{(1-p)\log_{p}t}{t}-\frac{\log_{p}(p+(1-p)t)}{p+(1-p)t}\right)\end{split}$
As to the right-hand side of (22), we have
@@
$G(t,p):=(1-p)(p+(1-p)t)\log_{p}t-p(1-p)\log_{p}(t)\cdot\log_{p}(p+(1-p)t)-t\cdot\log_{p}(p+(1-p)t)\geq 0.$
@@
$\begin{split}\frac{\partial}{\partial t}G(t,p)&=\frac{1}{\ln(p)}\left(\frac{p(1-p)}{t}+(1-p)^{2}\ln(t)+(1-p)^{2}-\frac{p(1-p)}{\ln(p)}\cdot\frac{\ln(p+(1-p)t)}{t}\right)\\ &-\frac{1}{\ln(p)}\left(\frac{p(1-p)^{2}}{\ln(p)}\cdot\frac{\ln(t)}{p+(1-p)t}+\ln(p+(1-p)t)+\frac{(1-p)t}{p+(1-p)t}\right).\end{split}$
@@
$\begin{split}&\frac{\partial}{\partial t}H(t,p)=\frac{1}{\ln(p)^{2}}\left[-(p-1)^{2}p\left(2+\ln(t)+\ln(p+t-pt)\right)\right]\\ &+\frac{1}{\ln(p)}\left[(-2t+p(2t-1)\ln(p+t-pt)\right]\\ &-\frac{1}{\ln(p)}\left[3(p-1)p(1+p(t-1)-2t)+(p-1)^{2}(-2t+p(2t-1))\ln(t)\right];\\ &\frac{\partial^{2}}{\partial t^{2}}H(t,p)=\frac{1}{\ln(p)^{2}}\left(-(p-1)^{2}p\left(\frac{1}{t}+\frac{1-p}{p+t-pt}\right)\right)\\ &+\frac{1}{\ln(p)}\left(-3(-2+p)(p-1)p-\frac{(p-1)^{2}(-2t+p(2t-1))}{t}+\frac{(1-p)(-2t+p(2t-1))}{p+t-pt}\right)\\ &+\frac{1}{\ln(p)}\left(-2(p-1)^{3}\ln(t)+2(p-1)\ln(p+t-pt)\right);\\ &\frac{\partial^{3}}{\partial t^{3}}H(t,p)=-\frac{(p-1)^{2}p}{t^{2}(p+t-pt)^{2}\ln(p)^{2}}\cdot h(t,p),\end{split}$

### v3

@@ 3.2 Induction Part
  • We may assume without loss of generality that $g\geq 0$, since replacing $g$ by its absolute value preserves the right-hand side of Eq. 8, while it only decreases the left-hand side:
@@
  • We may assume the support of $g$ is the whole set $A$. Indeed, if we denote $B:={\rm supp}(g)=\{x\in\{0,1\}^{n}:g(x)>0\}$ with $B\subset A$. Then $\mathbb{E}_{p}[g]^{2}=\left(\sum_{x\in B}g(x)\mu_{p}(x)\right)^{2}$. Hence, replacing the support of $g$ with the whole set $A$ only increases the right-hand side of Eq. 8.
@@
$p\mathcal{E}_{p}^{n-1}(g_{1},g_{1})\geq\frac{\mathbb{E}_{p}[g_{1}]^{2}}{a_{1}}\log_{p}a_{1},\quad p\mathcal{E}_{p}^{n-1}(g_{0},g_{0})\geq\frac{\mathbb{E}_{p}[g_{0}]^{2}}{a_{0}}\log_{p}a_{0}.$ (30)
@@
    where the first inequality is due to $a_{j}=\mu_{p}^{n-1}(A_{j})\leq 1$ and the second inequality is due to Jensen’s inequality. Hence replacing $g_{j}$ with their averages on the corresponding subsets can only decrease the left-hand side and does not change the right-hand side in the second inequality of Eq. 31.
@@
$\begin{split}&s_{1}:=\mathbb{E}_{p}[g_{1}]=\sum\limits_{x\in\{0,1\}^{n-1}}\mathbbm{1}_{A_{1}}(x)\cdot g_{1}(x)\mu_{p}^{n-1}(x)\\ &\quad\ =\sum\limits_{x\in A_{0}}\mathbbm{1}_{A_{1}}(x)\cdot g_{1}(x)\mu_{p}^{n-1}(x)+\sum\limits_{x\in A_{1}\setminus A_{0}}\mathbbm{1}_{A_{1}}(x)\cdot g_{1}(x)\mu_{p}^{n-1}(x)\\ &\quad\ =\alpha a_{0}+\gamma(a_{1}-a_{0}),\\ &s_{0}:=\mathbb{E}_{p}[g_{0}]=\sum\limits_{x\in\{0,1\}^{n-1}}\mathbbm{1}_{A_{0}}(x)\cdot g_{0}(x)\mu_{p}^{n-1}(x)\\ &\quad\ =\sum\limits_{x\in A_{0}}\mathbbm{1}_{A_{0}}(x)\cdot g_{0}(x)\mu_{p}^{n-1}(x)+\sum\limits_{x\in A_{1}\setminus A_{0}}\mathbbm{1}_{A_{0}}(x)\cdot g_{0}(x)\mu_{p}^{n-1}(x)=\beta a_{0},\\ &\|g_{1}-g_{0}\|_{2,\mu_{p}}^{2}=\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}[(g_{1}(x^{\prime})-g_{0}(x^{\prime}))^{2}]\\ &\quad\ =\sum\limits_{x^{\prime}\in A_{0}}(g_{1}(x^{\prime})-g_{0}(x^{\prime}))^{2}\mu_{p}^{n-1}(x^{\prime})+\sum\limits_{x^{\prime}\in A_{1}\setminus A_{0}}(g_{1}(x^{\prime})-g_{0}(x^{\prime}))^{2}\mu_{p}^{n-1}(x^{\prime})\\ &\quad\ =(\alpha-\beta)^{2}\cdot a_{0}+(a_{1}-a_{0})\cdot\gamma^{2}.\end{split}$
@@
$\begin{split}ps_{1}^{2}f(a_{1})+(1-p)s_{0}^{2}f(a_{0})&+pa_{0}(\alpha-\beta)^{2}+p(a_{1}-a_{0})\gamma^{2}\\ &\geq(ps_{1}+(1-p)s_{0})^{2}f(pa_{1}+(1-p)a_{0})).\end{split}$ (33)
@@ 3.3 Two-point Inequality Part
Renaming the variables $x=a_{0}$ and $y=a_{1}$, and recalling the constraints on $a_{0}$ and $a_{1}$, we need to prove Eq. 35 for $0<x<y\leq 1$.
$\begin{split}&p\cdot f(y)+(1-p)\cdot f(x)+(1-p)\cdot yf(x)f(y)\\ &\geq(p(1-p)yf(x)+(1-p)^{2}yf(y)+1)\cdot f(py+(1-p)x).\end{split}$ (36)
@@
$\begin{split}&\Delta_{p}(x,y)=\Delta_{p}(ty,y)=pf(y)+(1-p)f(t\cdot y)-f((p+(1-p)t)\cdot y)\\ &=pf(y)+\frac{1}{t}\cdot(1-p)f(y)+\frac{1}{y}\cdot(1-p)f(t)-\frac{f(y)}{p+(1-p)t}-\frac{f(p+(1-p)t)}{y}\\ &=\frac{p(1-p)(t-1)^{2}}{t(p+(1-p)t)}\cdot f(y)+\frac{1}{y}\cdot\left(\frac{(1-p)\log_{p}t}{t}-\frac{\log_{p}(p+(1-p)t)}{p+(1-p)t}\right).\end{split}$
As to the right-hand side of Eq. 37, we have
@@
$\begin{split}&G(t,p):=(1-p)(p+(1-p)t)\log_{p}t\\ &-p(1-p)\log_{p}(t)\cdot\log_{p}(p+(1-p)t)-t\cdot\log_{p}(p+(1-p)t)\geq 0.\end{split}$
@@
$\begin{split}\frac{\partial}{\partial t}G(t,p)&=\frac{1}{\ln(p)}\left(\frac{p(1-p)}{t}+(1-p)^{2}\ln(t)+(1-p)^{2}\right)\\ &-\frac{1}{\ln(p)}\left(\frac{p(1-p)}{\ln(p)}\cdot\frac{\ln(p+(1-p)t)}{t}\right)\\ &-\frac{1}{\ln(p)}\left(\frac{p(1-p)^{2}}{\ln(p)}\cdot\frac{\ln(t)}{p+(1-p)t}+\ln(p+(1-p)t)+\frac{(1-p)t}{p+(1-p)t}\right).\end{split}$
@@
$\begin{split}&\frac{\partial}{\partial t}H(t,p)=\frac{1}{\ln(p)^{2}}\left[-(p-1)^{2}p\left(2+\ln(t)+\ln(p+t-pt)\right)\right]\\ &+\frac{1}{\ln(p)}\left[(-2t+p(2t-1)\ln(p+t-pt)\right]\\ &-\frac{1}{\ln(p)}\left[3(p-1)p(1+p(t-1)-2t)+(p-1)^{2}(-2t+p(2t-1))\ln(t)\right];\\ &\frac{\partial^{2}}{\partial t^{2}}H(t,p)=\frac{1}{\ln(p)^{2}}\left(-(p-1)^{2}p\left(\frac{1}{t}+\frac{1-p}{p+t-pt}\right)\right)\\ &+\frac{1}{\ln(p)}\left(-3(-2+p)(p-1)p-\frac{(p-1)^{2}(-2t+p(2t-1))}{t}\right)\\ &+\frac{1}{\ln(p)}\left(\frac{(1-p)(-2t+p(2t-1))}{p+t-pt}-2(p-1)^{3}\ln(t)+2(p-1)\ln(p+t-pt)\right);\\ &\frac{\partial^{3}}{\partial t^{3}}H(t,p)=-\frac{(p-1)^{2}p}{t^{2}(p+t-pt)^{2}\ln(p)^{2}}\cdot h(t,p),\end{split}$